OpenStudy (anonymous):

log (base 3) (x+2) = 2?

OpenStudy (jdoe0001):

\(\bf \Large{ log_3(x+3)=2 \\ \quad \\ \textit{log cancellation rule of }{\color{blue}{ a^{log_ax}=x}}\qquad thus \\ \quad \\ log_3(x+3)=2\implies {\color{blue}{ 3}}^{log_{\color{blue}{ 3}}(x+3)}={\color{blue}{ 3}}^2\\ \quad \\\implies x+3=3^2 }\)

OpenStudy (jdoe0001):

and I think you can take it from there